# Enhancing Performance of Explainable AI Models with Constrained Concept Refinement

### v1

@@ 1 Introduction
ML algorithms are often caught in a dilemma between interpretability and performance. Models such as linear regression [24] and decision trees [43] provide straightforward interpretability through parameter weights and rule-based predictions. However, they frequently underperform on complex tasks. On the other hand, high-performing models, such as deep neural networks [31] and large language models [59], are notoriously opaque, given their large parameter spaces and intricate architectures.
Numerous methods have been proposed to extract interpretability from complex models [10, 56, 67, 49, 48, 55, 29, 52]. However, these approaches typically adopt post-hoc strategies, utilizing sensitivity analysis to identify the key parameters that influence predictions. Consequently, these methods lack guarantees of providing explanations for random prediction or ensuring their trustworthiness [4, 46, 27, 21, 50].
An alternative solution is to develop models that are interpretable by design. These models intrinsically integrate transparency into their architecture, providing explanations directly tied to their predictions. Concept Bottleneck Models (CBMs, [30]) exemplify this approach by mapping input data to an intermediate representation of human-defined concepts, which is then used for prediction. While concept-based models are promising, they require datasets annotated with concept scores, limiting their applicability. Extensions of CBMs aim to improve the approach by utilizing pretrained encoders like CLIP [65, 40]. A recent line of research [15, 11, 14] has introduced a decision-tree-based architecture with enhanced explainability. The predictions of the model are generated through a sequence of queries, each closely associated with human-specified concepts.
@@
A high level design paradigm of expainable AI is illustrated in Figure 1. Within these models, concept embeddings play a pivotal role: they not only influence prediction generation but also serve as the source of interpretability. However, recent studies have raised several questions against its reliability. These challenges are centered around its ambiguity [35, 61, 28, 38], fragility [18], and low accuracy [66, 13]. For embeddings generated with pretrained models, their encoders are reported to face issues such as domain adaptation [54, 22, 19, 63], biases in pretraining data [25], and sensitivity to encoding errors [47].
Efforts to address these challenges include introducing unsupervised modules to enhance concept expressiveness [51], training residual learning modules in parallel with concept encoders [65], learning dual embeddings per concept [66], and adding specialized data encoders to address domain shifts [13]. However, these approaches all involve introducing additional black-box learning modules, which undermines the purpose of explainable AI. Moreover, theoretical guarantees for such approaches, which is of particular interest for safety-critical domains, such as medical imaging [8] and autonomous vehicles [6], remain largely unexplored.
@@
  • Theoretical necessity. We rigorously demonstrate the necessity of refining concept embeddings by establishing a non-vanishing worst-case lower bound on model performance when such refinements are not applied to the concepts (Theorem 2). This highlights and motivates the idea of concept embedding refinement, which is at the crux of our proposed method.
  • Accuracy and interpretability guarantees. We show that CCR overcomes the aforementioned challenge by eliminating performance degradation through the appropriate refinement of concepts (Theorem 3). Furthermore, to quantify the performance of CCR, we consider a generative model as a test-bed, where both “interpretability” and “accuracy” admit crisp mathematical definitions. Under this model, we demonstrate that CCR achieves zero training loss while progressively enhancing interpretability (Theorem 4).
  • Application in interpretable image classification. We demonstrate the practical efficacy of CCR on multiple image classification benchmarks including CIFAR 10/100 [26], ImageNet [16], CUB200 [62] and Places365 [68]. In all benchmarks except CUB-200, CCR outperforms two recently developed explainable methods for visual classification tasks—CLIP-IP-OMP [14] and label-free CBM (lf-CBM) [40]—in terms of prediction accuracy while preserving interpretability, achieving this with a tenfold reduction in runtime.
@@ 2 Revisiting IP-OMP
We consider a classical interpretable-by-design setting in the literature [15, 14], namely the task of predicting a random variable by sequentially selecting a list of related random variables, termed queries, and checking their values. The problem is defined as follows:
Definition 1 (Single Variable Prediction by Query Selection).
@@
Definition 1 encompasses a broad spectrum of ML problems with applications in animal identification [32], medical diagnosis [20, 12], visual question-answering [64], image retrieval [7], and so on.
@@
In [23], the authors introduced a classic greedy-style algorithm termed Information Pursuit (IP) to address the task outlined in Definition 1. Specifically, IP iteratively selects the most informative query based on the observed values of previously selected queries. We provide a formal definition of IP for subsequent discussion:
Definition 2 (Information Pursuit).
@@
In [11], IP is demonstrated to achieve highly interpretable predictions with competitive accuracies. However, a significant limitation of IP is its substantial computational expense, as mutual information depends not only on the set $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ but also on previous observations, leading to an exponentially large input space. [14] addresses this complexity by imposing additional assumptions on the underlying generative model. The first quantifies the connection between $y$ and $q_{i}$, while the second defines their connections with ${\mathbf{d}}_{i}$.
Assumption 1 (Generative model for IP-OMP).
@@
Assumption 2 (Prior knowledge of latent feature vectors).
@@
Under 1 and 2, [14] establish a connection between IP and Orthogonal Matching Pursuit (OMP) [42], a seminal algorithm in the field of sparse coding.
Theorem 1 (IP-OMP [14]).
Under 1 and 2 and upon setting ${\mathbf{d}}_{i}={\mathbf{v}}_{i}$ for $i=1,\dots,n$, the selection rule $\pi(\cdot)$ defined in Definition 2 admits a closed-form expression given by:
@@
The main theoretical contribution of Theorem 1 lies in its ability to transform the computation of mutual information into that of a linear projection, thereby significantly reducing computational complexity. Although the equivalence is derived under simplifying assumptions, IP-OMP demonstrates remarkable generalization capabilities across various benchmark experiments, outperforming models based on the classical IP algorithm [14].
@@ 2.1 A Major Hurdle: Inaccurate Query Features
As discussed in the previous section, Theorem 1, and as a result, the success of IP-OMP is contingent upon correctly choosing the query features $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$.
For learning algorithms, the effective utilization of $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ is crucial for the accurate prediction of $y$. On the other hand, humans derive comprehension from $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ as these elements are intended to embed concepts that are interpretable by humans. In practical applications, $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ are either embedded and learned from predefined datasets where concepts are explicitly labeled [30] or generated by pretrained multimodal models, with CLIP being the most popular example [40].
However, these learned embeddings are often misaligned or inaccurate due to the inherent ambiguity and noise in the training of the models used to generate. For example, unlike carefully curated datasets, CLIP relies on large-scale, noisy image-text pairs scraped from the internet. These pairs often include mislabeled data, vague descriptions, or culturally biased associations, resulting in inconsistencies in representation, as recently reported by [17].
This suggests that 2 may be overly optimistic. In this section, we will illustrate how the violation of 2—specifically, deviations of the available query features from the latent feature vectors—results in a proportional degradation in the performance of IP-OMP. For the remainder of this paper, we denote the observed query feature set compactly by the matrix ${\bm{D}}=\begin{bmatrix}{\mathbf{d}}_{1}&{\mathbf{d}}_{2}&\cdots&{\mathbf{d}}_{n}\end{bmatrix}\in\mathbb{R}^{d\times n}$. This matrix ${\bm{D}}$ is also referred to as a feature query matrix. We also refer to ${\bm{D}}^{*}=\begin{bmatrix}{\mathbf{v}}_{1}&{\mathbf{v}}_{2}&\cdots&{\mathbf{v}}_{n}\end{bmatrix}\in\mathbb{R}^{d\times n}$ as the ground truth latent feature matrix. 2 can now be interpreted as assuming ${\bm{D}}={\bm{D}}^{*}$, stating that we are using the ground truth latent feature vectors as the feature query matrix. This matrix can then be utilized to determine which queries should be observed.
Let $f_{{\bm{D}}}({\mathbf{x}},\{r_{\pi(i)}\}_{i=1}^{k})$ be the maximum likelihood estimator of $y$, defined as Definition 2, when the selection rule $\pi(\cdot)$ is obtained by executing IP-OMP using ${\bm{D}}$ as the query feature set in Theorem 1. For brevity, when the context is clear, we refer to this maximum likelihood estimator simply as $f_{{\bm{D}}}$. Moreover, define $\mathcal{L}(y_{\mathrm{pred}})$ as the squared population loss of any estimator $y_{\mathrm{pred}}$:
@@
where ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$ is the unknown random vector used in the generative model described in 1.
To showcase the effect of deviation in the query features, we consider a scenario where the input ${\mathbf{x}}$ can be written as a linear combination of $k$ latent feature vectors (corresponding to $k$ columns of ${\bm{D}}^{*}$). This assumption is prevalent in the sparse coding literature [5, 2, 34] and ensures that using ${\bm{D}}^{*}$ as the query feature set guarantees good prediction performance. However, even in this ideal scenario, we demonstrate that even a small column-wise deviation from ${\bm{D}}^{*}$ can lead to performance degradation of IP-OMP.
Theorem 2.
Suppose that 1 holds. Furthermore, suppose that ${\bm{D}}^{*}$ is column-orthogonal and ${\mathbf{x}}={\bm{D}}^{*}\mathbf{\beta}$ for some $k$-sparse vector $\mathbf{\beta}$ with $\|\mathbf{\beta}\|_{0}=k$. Additionally, assume that the non-zero entries in $\mathbf{\beta}$ have absolute values bounded below by $\gamma$ and above by $\Gamma$. Then, for any $\epsilon\in\left(0,\frac{1}{\sqrt{1+16\Gamma^{2}/\gamma^{2}}}\right)$, there exists another orthonormal matrix $\tilde{{\bm{D}}}\in\mathbb{R}^{d\times n}$ such that $\|\tilde{{\bm{D}}}-{\bm{D}}^{*}\|_{1,2}\leq\epsilon$ and
@@
We present the complete proof of Theorem 2 in Section C.1, which proceeds as follows: (1) we derive the closed form solution for $\mathcal{L}(f_{{\bm{D}}})$ using the column-orthogonality of ${\bm{D}}$ and the distribution of ${\mathbf{z}}$; (2) we then construct the example that satisfies Equation 2 by rotating columns of ${\bm{D}}^{*}$ alongside the subspace spanned by itself.
When $\Gamma/\gamma$ is bounded by a constant, we have $\|{\mathbf{x}}\|^{2}_{2}=\|\mathbf{\beta}\|^{2}_{2}=\mathbf{\Theta}(k\gamma^{2})$ which simplifies Equation 2 to $\mathcal{L}(f_{\tilde{{\bm{D}}}})-\mathcal{L}(f_{{\bm{D}}^{*}})=\Omega(\epsilon^{2}\|{\mathbf{x}}\|^{2}_{2})$. This indicates that the perturbation $\epsilon$ is fully captured by the resulting gap in the squared loss, scaled by a factor of $\|{\mathbf{x}}\|^{2}_{2}$.

### v2

@@ 1 Introduction
ML algorithms are often caught in a dilemma between interpretability and performance. Models such as linear regression (Hastie, 2009) and decision trees (Quinlan, 1986) provide straightforward interpretability through parameter weights and rule-based predictions. However, they frequently underperform on complex tasks. On the other hand, high-performing models, such as deep neural networks (LeCun et al., 2015) and large language models (Vaswani, 2017), are notoriously opaque, given their large parameter spaces and intricate architectures.
Numerous methods have been proposed to extract interpretability from complex models (Baehrens et al., 2010; Simonyan et al., 2013; Zeiler & Fergus, 2014; Shrikumar et al., 2017; Selvaraju et al., 2017; Smilkov et al., 2017; Kolek et al., 2020; Subramanya et al., 2019). However, these approaches typically adopt post-hoc strategies, utilizing sensitivity analysis to identify the key parameters that influence predictions. Consequently, these methods lack guarantees of providing explanations for random prediction or ensuring their trustworthiness (Adebayo et al., 2018; Rudin, 2019; Kindermans et al., 2019; Ghorbani et al., 2019; Slack et al., 2020).
An alternative solution is to develop models that are interpretable by design. These models intrinsically integrate transparency into their architecture, providing explanations directly tied to their predictions. Concept Bottleneck Models (CBMs, (Koh et al., 2020)) exemplify this approach by mapping input data to an intermediate representation of human-defined concepts, which is then used for prediction. While concept-based models are promising, they require datasets annotated with concept scores, limiting their applicability. Extensions of CBMs aim to improve the approach by utilizing pretrained encoders like CLIP (Yuksekgonul et al., 2022; Oikarinen et al., 2023). A recent line of research (Chattopadhyay et al., 2022, 2023, 2024) has introduced a decision-tree-based architecture with enhanced explainability. The predictions of the model are generated through a sequence of queries, each closely associated with human-specified concepts.
@@
A high level design paradigm of expainable AI is illustrated in Figure 1. Within these models, concept embeddings play a pivotal role: they not only influence prediction generation but also serve as the source of interpretability. However, recent studies have raised several questions against its reliability. These challenges are centered around its ambiguity (Margeloiu et al., 2021; Watson, 2022; Kim et al., 2023; Marconato et al., 2023), fragility (Furby et al., 2024), and low accuracy (Zarlenga et al., 2022; Chowdhury et al., 2024). For embeddings generated with pretrained models, their encoders are reported to face issues such as domain adaptation (Shao et al., 2023; Gondal et al., 2024; Feng et al., 2024; Wang & Kang, 2024), biases in pretraining data (Hamidieh et al., 2024), and sensitivity to encoding errors (Ranjan et al., 2024).
Efforts to address these challenges include introducing unsupervised modules to enhance concept expressiveness (Sawada & Nakamura, 2022), training residual learning modules in parallel with concept encoders (Yuksekgonul et al., 2022), learning dual embeddings per concept (Zarlenga et al., 2022), and adding specialized data encoders to address domain shifts (Chowdhury et al., 2024). However, these approaches all involve introducing additional black-box learning modules, which undermines the purpose of explainable AI. Moreover, theoretical guarantees for such approaches, which is of particular interest for safety-critical domains, such as medical imaging (Barragán-Montero et al., 2021) and autonomous vehicles (Bensalem et al., 2023), remain largely unexplored.
@@
  • Theoretical necessity. We rigorously demonstrate the necessity of refining concept embeddings by establishing a non-vanishing worst-case lower bound on model performance when such refinements are not applied to the concepts (Theorem 2.6). This highlights and motivates the idea of concept embedding refinement, which is at the crux of our proposed method.
  • Accuracy and interpretability guarantees. We show that CCR overcomes the aforementioned challenge by eliminating performance degradation through the appropriate refinement of concepts (Theorem 3.3). Furthermore, to quantify the performance of CCR, we consider a generative model as a test-bed, where both “interpretability” and “accuracy” admit crisp mathematical definitions. Under this model, we demonstrate that CCR achieves zero training loss while progressively enhancing interpretability (Theorem 3.4).
  • Application in interpretable image classification. We demonstrate the practical efficacy of CCR on multiple image classification benchmarks including CIFAR 10/100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009), CUB200 (Wah et al., 2011) and Places365 (Zhou et al., 2017). In all benchmarks except CUB-200, CCR outperforms two recently developed explainable methods for visual classification tasks—CLIP-IP-OMP (Chattopadhyay et al., 2024) and label-free CBM (lf-CBM) (Oikarinen et al., 2023)—in terms of prediction accuracy while preserving interpretability, achieving this with a tenfold reduction in runtime.
On a high level, our main contribution is the introduction of a framework in which concept embeddings in explainable AI models are refined within a restricted parameter space to attain a better balance between predictive performance and interpretability. The remainder of the paper substantiates this claim along two key dimensions:
  1. 1.
    Theoretical validation. We demonstrate that our method is both necessary (Section 2.1) and effective (Section 3) under the theoretical framework introduced by (Chattopadhyay et al., 2024), whose background is reviewed in Section 2. This framework serves as a suitable testbed for our contributions for two main reasons: (1) Among interpretable-by-design models (Zhou et al., 2018; Koh et al., 2020; Yuksekgonul et al., 2022; Oikarinen et al., 2023; Chattopadhyay et al., 2023), only (Chattopadhyay et al., 2024) presents a generative model wherein both performance and interpretability are rigorously defined. (2) The algorithm motivated by this generative framework achieves state-of-the-art results within the class of interpretable AI methods.
  2. 2.
    Empirical evaluation. We conduct experiments on multiple benchmark datasets for image classification tasks to assess the practical effectiveness of our approach (Section 4).
@@ 2 Revisiting IP-OMP
We consider a classical interpretable-by-design setting in the literature (Chattopadhyay et al., 2022, 2024), namely the task of predicting a random variable by sequentially selecting a list of related random variables, termed queries, and checking their values. The problem is defined as follows:
Definition 2.1 (Single Variable Prediction by Query Selection).
@@
Definition 2.1 encompasses a broad spectrum of ML problems with applications in animal identification (Lampert et al., 2009), medical diagnosis (Graziani et al., 2018; Clough et al., 2019), visual question-answering (Yi et al., 2018), image retrieval (Bucher et al., 2019), and so on.
@@
In (Geman & Jedynak, 1996), the authors introduced a classic greedy-style algorithm termed Information Pursuit (IP) to address the task outlined in Definition 2.1. Specifically, IP iteratively selects the most informative query based on the observed values of previously selected queries. We provide a formal definition of IP for subsequent discussion:
Definition 2.2 (Information Pursuit).
@@
In (Chattopadhyay et al., 2023), IP is demonstrated to achieve highly interpretable predictions with competitive accuracies. However, a significant limitation of IP is its substantial computational expense, as mutual information depends not only on the set $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ but also on previous observations, leading to an exponentially large input space. (Chattopadhyay et al., 2024) addresses this complexity by imposing additional assumptions on the underlying generative model. The first quantifies the connection between $y$ and $q_{i}$, while the second defines their connections with ${\mathbf{d}}_{i}$.
Assumption 2.3 (Generative model for IP-OMP).
@@
Assumption 2.4 (Prior knowledge of latent feature vectors).
@@
Under 2.3 and 2.4, (Chattopadhyay et al., 2024) establish a connection between IP and Orthogonal Matching Pursuit (OMP) (Pati et al., 1993), a seminal algorithm in the field of sparse coding.
Theorem 2.5 (IP-OMP (Chattopadhyay et al., 2024)).
Under 2.3 and 2.4 and upon setting ${\mathbf{d}}_{i}={\mathbf{v}}_{i}$ for $i=1,\dots,n$, the selection rule $\pi(\cdot)$ defined in Definition 2.2 admits a closed-form expression given by:
@@
The main theoretical contribution of Theorem 2.5 lies in its ability to transform the computation of mutual information into that of a linear projection, thereby significantly reducing computational complexity. Although the equivalence is derived under simplifying assumptions, IP-OMP demonstrates remarkable generalization capabilities across various benchmark experiments, outperforming models based on the classical IP algorithm (Chattopadhyay et al., 2024).
@@ 2.1 A Major Hurdle: Inaccurate Query Features
As discussed in the previous section, Theorem 2.5, and as a result, the success of IP-OMP is contingent upon correctly choosing the query features $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$.
For learning algorithms, the effective utilization of $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ is crucial for the accurate prediction of $y$. On the other hand, humans derive comprehension from $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ as these elements are intended to embed concepts that are interpretable by humans. In practical applications, $\{{\mathbf{d}}_{i}\}_{i=1}^{n}$ are either embedded and learned from predefined datasets where concepts are explicitly labeled (Koh et al., 2020) or generated by pretrained multimodal models, with CLIP being the most popular example (Oikarinen et al., 2023).
However, these learned embeddings are often misaligned or inaccurate due to the inherent ambiguity and noise in the training of the models used to generate. For example, unlike carefully curated datasets, CLIP relies on large-scale, noisy image-text pairs scraped from the internet. These pairs often include mislabeled data, vague descriptions, or culturally biased associations, resulting in inconsistencies in representation, as recently reported by (Dutta et al., 2023).
This suggests that 2.4 may be overly optimistic. In this section, we will illustrate how the violation of 2.4—specifically, deviations of the available query features from the latent feature vectors—results in a proportional degradation in the performance of IP-OMP. For the remainder of this paper, we denote the observed query feature set compactly by the matrix ${\bm{D}}=\begin{bmatrix}{\mathbf{d}}_{1}&{\mathbf{d}}_{2}&\cdots&{\mathbf{d}}_{n}\end{bmatrix}\in\mathbb{R}^{d\times n}$. This matrix ${\bm{D}}$ is also referred to as a feature query matrix. We also refer to ${\bm{D}}^{*}=\begin{bmatrix}{\mathbf{v}}_{1}&{\mathbf{v}}_{2}&\cdots&{\mathbf{v}}_{n}\end{bmatrix}\in\mathbb{R}^{d\times n}$ as the ground truth latent feature matrix. 2.4 can now be interpreted as assuming ${\bm{D}}={\bm{D}}^{*}$, stating that we are using the ground truth latent feature vectors as the feature query matrix. This matrix can then be utilized to determine which queries should be observed.
Let $f_{{\bm{D}}}({\mathbf{x}},\{r_{\pi(i)}\}_{i=1}^{k})$ be the maximum likelihood estimator of $y$, defined as Definition 2.2, when the selection rule $\pi(\cdot)$ is obtained by executing IP-OMP using ${\bm{D}}$ as the query feature set in Theorem 2.5. For brevity, when the context is clear, we refer to this maximum likelihood estimator simply as $f_{{\bm{D}}}$. Moreover, define $\mathcal{L}(y_{\mathrm{pred}})$ as the squared population loss of any estimator $y_{\mathrm{pred}}$:
@@
where ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$ is the unknown random vector used in the generative model described in 2.3.
To showcase the effect of deviation in the query features, we consider a scenario where the input ${\mathbf{x}}$ can be written as a linear combination of $k$ latent feature vectors (corresponding to $k$ columns of ${\bm{D}}^{*}$). This assumption is prevalent in the sparse coding literature (Arora et al., 2015; Agarwal et al., 2016; Liang et al., 2022) and ensures that using ${\bm{D}}^{*}$ as the query feature set guarantees good prediction performance. However, even in this ideal scenario, we demonstrate that even a small column-wise deviation from ${\bm{D}}^{*}$ can lead to performance degradation of IP-OMP.
Theorem 2.6.
Suppose that 2.3 holds. Furthermore, suppose that ${\bm{D}}^{*}$ is column-orthogonal and ${\mathbf{x}}={\bm{D}}^{*}\bm{\beta}$ for some $k$-sparse vector $\bm{\beta}$ with $\|\bm{\beta}\|_{0}=k$. Additionally, assume that the non-zero entries in $\bm{\beta}$ have absolute values bounded below by $\gamma$ and above by $\Gamma$. Then, for any $\epsilon\in\left(0,\frac{1}{\sqrt{1+16\Gamma^{2}/\gamma^{2}}}\right)$, there exists another orthonormal matrix $\tilde{{\bm{D}}}\in\mathbb{R}^{d\times n}$ such that $\|\tilde{{\bm{D}}}-{\bm{D}}^{*}\|_{1,2}\leq\epsilon$ and
@@
We present the complete proof of Theorem 2.6 in Section C.1, which proceeds as follows: (1) we derive the closed form solution for $\mathcal{L}(f_{{\bm{D}}})$ using the column-orthogonality of ${\bm{D}}$ and the distribution of ${\mathbf{z}}$; (2) we then construct the example that satisfies Equation 2 by rotating columns of ${\bm{D}}^{*}$ alongside the subspace spanned by itself.
When $\Gamma/\gamma$ is bounded by a constant, we have $\|{\mathbf{x}}\|^{2}_{2}=\|\bm{\beta}\|^{2}_{2}=\mathbf{\Theta}(k\gamma^{2})$ which simplifies Equation 2 to $\mathcal{L}(f_{\tilde{{\bm{D}}}})-\mathcal{L}(f_{{\bm{D}}^{*}})=\Omega(\epsilon^{2}\|{\mathbf{x}}\|^{2}_{2})$. This indicates that the perturbation $\epsilon$ is fully captured by the resulting gap in the squared loss, scaled by a factor of $\|{\mathbf{x}}\|^{2}_{2}$.